OpenStudy (anonymous):

Given the functions f(x) = 3x - 1 and g(x) = 3x + 4, which operation results in the smallest coefficient on the x term?

OpenStudy (anonymous):

h(x) = x + 5, work alone h(x) = x + 5, team with Gertrude h(x) = 9x + 35, work alone h(x) = 9x + 35, team with Gertrude